OpenStudy (sleepyjess):

Find the standard form of the equation of the parabola with a vertex at the origin and a focus at (0, 9).

OpenStudy (sleepyjess):

@ganeshie8

OpenStudy (misty1212):

HI!!

OpenStudy (mathmate):

 For the parabola above, the equation is y=x^2/(4p) where p=distance from vertex to focus=distance from vertex to directrix. Since you know the vertex (0,0), and the focus (0,9), you should be able to find p, and hence the equation of the parabola.

OpenStudy (misty1212):

do you know what it looks like? if so, this takes one second

OpenStudy (sleepyjess):

p = 9

OpenStudy (misty1212):

yes of course

OpenStudy (misty1212):

making this \[36y=x^2\]

OpenStudy (sleepyjess):

Would that be equal to y = \(\dfrac1{36}x^2\)?

OpenStudy (sleepyjess):

Or \(y = \dfrac{1}{36}x^2\)

OpenStudy (sleepyjess):

@misty1212

OpenStudy (misty1212):

they are the same right?

OpenStudy (misty1212):

usually it is \[4p(y-k)=(x-h)^2\] as a standard form

OpenStudy (misty1212):

but \[36y=x^2\] is the same as \[y=\frac{1}{36}x^2\]

OpenStudy (sleepyjess):

Thank you so much! :)